# Blockchain-Enabled Privacy-Preserving Second-Order Federated Edge Learning in Personalized Healthcare

### v1

@@ III-D Communication Model
The system employs a layered communication model and protocol architecture designed to ensure privacy-preserving, secure, and auditable operations across asynchronous phases. Data collection is initiated by sensors and edge client devices, which acquire raw physiological metrics such as blood glucose levels and electrocardiogram signals. HTTPS/gRPC encrypts federated learning client-server communications, while MQTT optimizes lightweight data transfer from IoT/wearables to the edge client servers.
@@
Our proposed access scheme is used along with the widely used Elliptic Curve Integrated Encryption Scheme (ECIES), Child Key Derivation function (CKD), and the Elliptic Curve Digital Signature Algorithm (ECDSA) are cryptographic techniques to ensure secure data storage and communication. ECIES works by independently deriving a bulk encryption key and a MAC key from a common secret. The data is first encrypted under a symmetric cipher, and then the ciphertext is authenticated under an authentication scheme. Finally, the common secret is encrypted under the public part of a public/private key pair. The CKD function is used for managing data in batches; child key derivation functions are used in hierarchical deterministic wallets (HD wallets). It helps in generating a tree of keys from a single master key, which can be very useful for managing multiple keys securely and systematically. ECDSA is a digital signature algorithm that is used for secure key sharing for both data and communication. The ECDSA ensures that the data and communication are coming from the stated sender (authenticity), have not been altered in transit (integrity), and repudiation by the sender can be disputed (non-repudiation). Implementing ECIES, CKD, and ECDSA in this proposed system provides a robust framework that ensures secure data storage and communication. The ECIES offers a strong encryption scheme for data protection, the CKD provides an efficient way to manage data in batches, and the ECDSA guarantees secure key sharing and data authenticity.
@@
Following this, local model training at edge client level is completed using FedCurv, model local weights $F_{k}$, $g_{k}$ are transmitted through private ethereum network, ensuring the confidentiality of individual client $k$ data weights, preventing re-identification. Subsequently, processed data is transmitted to the model aggregation servers via a blockchain network, where aggregated global models are cryptographically hashed and immutably logged on the blockchain ledger. Transaction validation occurs through a distributed consensus protocol, ensuring integrity. Central to access governance is a patient-centric framework: patients configure data sharing permissions via a portal, triggering Ethereum based smart contracts, while researchers and clinicians submit queries via blockchain transactions, which undergo automated authorization checks. Approved requests retrieve models or insights from secure decentralized storage systems such as IPFS, with all access events (identity, timestamp, purpose) permanently recorded on-chain for auditability. Role-Based Access Control (RBAC) policies are programmatically enforced through smart contracts.
@@ IV Experimental setup
To evaluate the performance of our proposed BFEL framework through experimental testbeds, optimized FedCurv is employed as federated learning algorithm to address data heterogeneity across clients. FedCurv is compared with baseline federated algorithm FedAvg. The federated learning framework involves 10 local training rounds per participant and 20 global aggregation rounds. We utilise three benchmark datasets: MNIST, CIFAR10, and MedMNIST. MNIST and CIFAR10 were selected as standard benchmarks in neural network research, while MedMNIST specifically its PathMNIST subset, derived from 2D image classification. On the client side, each edge device performs local training while incorporating a curvature-aware regularization term. This regularization penalizes deviations from the global model based on the estimated importance of each parameter, thereby preserving critical knowledge and improving model stability. On the server side 2, the curvature matrices and gradients collected from the clients are aggregated. The server 1 then applies a curvature-scaled update to the global model, ensuring that parameter updates are inversely proportional to their estimated importance. This approach enables more efficient and robust federated optimization in non-iid settings. The private ethereum blockchain network is utilised as a service layer demonstrating its capability to uphold auditability, verifiability, and availability while ensuring privacy preservation in edge-FL settings.
@@ IV-A Parameter Settings
MNIST, a standard benchmark for image classification, includes 60,000 images (28×28 pixels), split into 50,000 training and 10,000 test samples, providing a baseline for foundational algorithm validation. MedMNIST subclass PathMNIST (28×28 pixels), a more challenging alternative, comprises 35,000 training images and 8000 testing images in non-iid, enabling evaluation of model generalizability in scenarios with higher intra-class variability. CIFAR-10, a widely used dataset for object recognition, offers 60,000 color images (32×32 pixels). Together, these datasets simulate real-world edge-FL challenges, such as decentralized computation and heterogeneous data privacy requirements, while validating the framework’s ability to balance transparency, security, and efficiency in privacy-sensitive environments.
@@ IV-B Network Configuration:
This study utilizes two widely recognized neural architectures: a multi-layer perceptron (MLP) and a convolutional neural network (CNN). These models served as the foundational deep learning frameworks for training classification systems within a federated learning setup, simulating server client training scenarios. The experiments aimed to assess how effectively each algorithm handles non-iid data and preserves knowledge across distributed medical and non-medical imaging tasks. For MNIST and PathMNIST (image datasets), the CNN consists of two convolutional layers, each followed by a max-pooling layer, and concludes with two fully connected layers. To accommodate the RGB input channels of the CIFAR-10 dataset, the architecture is modified while retaining the core structure. Both configurations employ the SGD optimizer with a learning rate of 0.001.
@@ IV-C Hardware and Software Configuration
We utilise raspberry Pi3 model B+ minicomputers as edge servers (manager nodes of side chains) and lightweight nodes are implemented using STM32F427 development boards (low-power ARM Cortex M3,M4 and M7 processors), which are used for high-speed implementation of asymmetric cryptographic algorithms. The elliptic curve digital signature algorithm (ECDSA) is used to generate public and private keys, and device authentication mechanisms. The STMicroelectronics X-CUBE-CRYPTOLIB library is utilized to implement several standard cryptographic algorithms with the ARM Cortex-M series processors. System components are developed using Go (Golang), solidity to write smart contracts and deployed using remix IDE, and MetaMask handles concurrent transactions and interactions with third-party cloud services. A PoS consensus ensures block validation, while the gossip protocol enables fast, resilient message propagation and node synchronization. Figure 8 depicts the system’s initialized nodes.
@@
         Compute regularized loss
@@ V Performance Evaluations
To validate the effectiveness of BFEL, we conducted experiments across varying task sequence configurations using FedCurv and compared its performance against established baselines using FedAvg. The results are analyzed through multiple perspectives, offering distinct insights. Code is available publically at our github account [^1] https://github.com/AnumNawazKahloon/FedCurv.
@@ V-A Federated Simulation Results
From the epochs per round standpoint, accuracy consistently improves as the number of local training epochs increases across all settings and algorithms. It depicts the close alignment of local optima with global optima, making extended local training within each round advantageous when maintaining a fixed number of communication rounds. Accuracy of FedAvg increases sharply after each epoch as compared to FedCurv at the base level depicted in figure 7 and  6 as well as edge client level, which shows the minimum divergence of results after each round in FedCurv, makes it more consistent and according to the previous weights of edge client device.
Overall comparison of performance, FedCurv designed primarily to address non-iid data challenges in federated learning, surprisingly outperforms FedAvg even in uniform data settings. Notably, FedCurv often achieves superior accuracy after 100 rounds, indicating a potential need for extended training phases to reach convergence compared to FedAvg. Regarding communication efficiency, reducing the frequency of communication rounds while keeping the total number of training epochs constant yields better model performance, implying that less frequent parameter exchanges may enhance learning stability or optimization. The experimental setup utilized an SGD optimizer with adaptive learning rate decay (reduced by a factor of 3 every 5 epochs if validation loss stagnated), a mini-batch size of 20, 20 rounds per task, and 1 epoch per round. For the MLP configuration, adjustments included a reduced mini-batch size of 10 and an initial learning rate of $1\mathrm{e}{-4}$, while client sampling fractions of 0.25 and 0.05 were applied at each round. Hyperparameters $\lambda 1$ and $\lambda 2$ were fixed at $[1\mathrm{e}{-1},4\mathrm{e}{-1}]$ and 100, respectively, across all experiments. These findings collectively highlight the interplay between local training intensity, algorithmic robustness, and communication strategies in federated learning frameworks.
@@ V-B Layer2 Ethereum Implementation
The second layer ethereum network is implemented using a polygon sidechain structure which leverages proof-of-stake (PoS) consensus. It consists of a main blockchain layer running on cloud servers and a sub-blockchain network comprised of sidechains. The main blockchain layer implements a consortium blockchain, while the sub-blockchain layer consists of multiple private sidechains, each deployed on individual edge servers to facilitate localized model training in BFEL. Within each sub-blockchain, the corresponding edge server functions as a miner node, responsible for performing local model training and updating local model weights based on the global model parameters after each training epoch. Subsequently, each edge server transmits its locally trained results to the main blockchain layer, where global model parameters are aggregated.
@@
The edge servers are also tasked with collecting and packaging model training outputs (treated as transactions) into candidate blocks, which are then published to their respective private blockchains. The main blockchain layer is operated by cloud servers configured as a consortium blockchain. Each registered user organization such as hospitals or governmental agencies within the consortium is permitted to establish its own private sidechain. Furthermore, to reduce potential latency caused by complex task assignment mechanisms, transaction packaging responsibilities in both private and consortium blockchains are delegated to designated manager nodes.
@@ V-C Security Analysis
We have conducted a performance evaluation of three different types of STM32F427 M series processors within the framework of asymmetric cryptography. We utilized the X-CUBE-CRYPTOLIB library to implement the ECDSA. To ascertain the statistical error of the results obtained over the number of executions, we calculated the mean, standard deviation, and standard error using the appropriate equations.
@@
Power consumption is an important parameter of microcontrollers, we determine power utilization during the execution of cryptographic algorithms. The power consumed was determined by measuring the voltage across a shunt resistance R. The current consumption of the processors was calculated using ohm’s law. To determine the average power consumption, ECDSA was executed for 15 runs and a comparison of power consumption is presented in figure 9(b). The average power consumption by M3 and M4 was $\pm$ 200mW, whereas M7 used an average of $\pm$ 290mW. Results indicate the superior performance of M4 cortex microcontrollers are best fit while consuming fewer resources.
@@
[…]
TABLE II: Security Measures
Parameter Implementation
Authorization Public key cryptography to encrypt CSK
Confidentiality Public key cryptography and proof of authority
Integrity Broadcast hash of each data_batch
Availability Achieved by limiting number of requests
Anonymity Discard raw data and store only processed information
@@ V-D Scalability Analysis
Edge clients based DLTs offer several benefits but also face inherent limitations in terms of scalability, which limit their application range. Nevertheless, BFEL handles this challenge using the side chains concept, gossip protocol, and PoS consensus. FoBSim simulation tool [^2] https://github.com/sed-szeged/FobSim is utilized to check the scalability of the proposed model. Manager nodes ranging from 5 to 500 were used to check the performance matrix of the proposed model and measure the total time required to complete the request procedure at edge clients versus at the cloud layer.
The edge client measures and divides the time needed to complete a transaction into three sections: Time to retrieve data (TRD), time for checking the transaction (VTR), and time for confirming the transaction (TCT). Figure 10(a) presents the results of the measurements. Based on the available resources, it is impressive that TRD requires only 34.6 milliseconds on average, VTR 36 milliseconds on average, and TCT 73.6 milliseconds on average. Additionally, it is essential to note that TCT also relies on the network, which in this experiment was adversely affected by our shared Wi-Fi’s slow response time, causing the time to be extended overall.
Gossip protocols are integrated to improve the scalability of the proposed network model and public key infrasture (PKI) is utilised for cryptographically signing model updates and transactions. This architecture integrates end-to-end privacy preservation, tamper evident logging, and granular access controls, ensuring regulatory compliance and transparency across all operational phases.
The impact of the gossip protocol on the total elapsed time at edge clients was evaluated to demonstrate its scalability enhancement. Figure 10(b) illustrates the improved scalability of the proposed access scheme when utilizing the gossip protocol, as compared to transactions conducted without it.
Figure 11(a) shows that the cloud layer utilises around double the time as compared to edge clients to complete transaction requests during concurrent transactions starting from 5 to 100 transactions at a time.
@@ VI Discussion and Analysis
In this study, we propose and develop a personalized healthcare system BFEL based on second order FEL. Second-order FL methods bring a substantial benefit of personalized training through their capability to use loss function curvature data to improve personalized training procedures during local data training. BFEL leverages blockchain as a service layer to ensure the privacy, transparency and confidentiality of data owners while maintaining auditability, verifiability, availability, and robust privacy protection in FEL environments. Upon model training completion, the global model and local parameters are securely stored on the blockchain and subsequently distributed to the federated learning edge clients and cloud server and vice versa in accordance with the established policy. Smart contracts are utilized to broadcast policies while providing automated confirmation and distribution operations.
Optimized FedCurv is used to minimize the number of communication rounds needed for personalized non-iid and heterogeneous distributed training across edge client devices to enhance personalized health monitoring and prediction. Performance evaluations demonstrate better performance of FedCurv and resistance capabilities than FedAvg using CNNs and MLPs in non-iid conditions. Key evaluation metrics including throughput, accuracy, privacy, and scalability analysis has been analyzed. Analysis of the experimental results demonstrates that the proposed method preserves the accuracy of the machine learning process while ensuring auditability and verifiability throughout the training and aggregation procedures. Privacy protection is achieved through second-order FEL using FedCurv and public key encryption ECDSA and ECIES, while latency and throughput were evaluated by measuring communication transactions on a permissioned blockchain and compared against a benchmark model FedAvg. The results demonstrate that the proposed model outperforms the benchmark by achieving enhanced privacy, accuracy and scalability. Results underscore better stability and accuracy of the proposed framework BFEL in heterogeneous settings and non-iid environments where patient data varies widely.

### v2

@@ III-D Communication Model
The system employs a layered communication model and protocol architecture designed to ensure privacy-preserving, secure, and auditable operations across asynchronous phases. Data collection is initiated by sensors and client devices, which acquire raw physiological metrics such as blood glucose levels and electrocardiogram signals. Https/gRPC encrypts FL client-server communications, while mqtt optimizes lightweight data transfer from IoT/wearables to the client servers.
@@
Our proposed access scheme is used along with the widely used elliptic curve integrated encryption scheme (ECIES), child key derivation function (ckd), and the elliptic curve digital signature algorithm (ECDSA) are cryptographic techniques to ensure secure data storage and communication. ECIES works by independently deriving a bulk encryption key and a mac key from a common secret. The data is first encrypted under a symmetric cipher, and then the ciphertext is authenticated under an authentication scheme. Finally, the common secret is encrypted under the public part of a public-private key pair. The ckd function is used for managing data in batches, child key derivation functions are used in hierarchical deterministic wallets (hd wallets). It helps in generating a tree of keys from a single master key, which can be very useful for managing multiple keys securely and systematically. ECDSA is a digital signature algorithm that is used for secure key sharing for both data and communication. The ECDSA ensures that the data and communication are coming from the stated sender (authenticity), have not been altered in transit (integrity), and repudiation by the sender can be disputed (non-repudiation). Implementing ECIES, ckd, and ECDSA in this proposed system pipeline provides a robust framework that ensures secure data storage and communication. The ECIES offers a strong encryption scheme for data protection, the ckd provides an efficient way to manage data in batches, and the ECDSA guarantees secure key sharing and data authenticity.
@@
Following this, local model training at client device level is completed using FedCurv, model local weights $F_{k}$, $g_{k}$ are transmitted through private ethereum network, ensuring the confidentiality of individual client $k$ data weights, preventing re-identification. Subsequently, processed data is transmitted to the model aggregation servers via a blockchain network, where aggregated global models are cryptographically hashed and immutably logged on the blockchain ledger. Transaction validation occurs through a distributed consensus protocol, ensuring integrity. In this patient-centric framework, patients configure data sharing permissions via a portal, triggering ethereum based smart contracts, while researchers and clinicians submit queries via blockchain transactions, which undergo automated authorization checks. Approved requests retrieve models or insights from secure decentralised storage systems such as IPFS, with all access events (identity, timestamp, purpose) permanently recorded on-chain for auditability. Role-based access control (RBAC) policies are programmatically enforced through smart contracts.
@@ IV Experimental setup
This section details the experimental setup designed to evaluate the performance of the proposed BFEL framework. To address data heterogeneity across clients, we employ the optimised FedCurv algorithm as our core federated learning method and compare its performance against the baseline algorithm, FedAvg. The FL network is configured with each participant performing 10 local training rounds per global round, and 20 global aggregation rounds. We utilise three benchmark datasets: mnist, cifar-10, and medmnist. mnist and cifar-10, as standard benchmarks in neural network research, while medmnist specifically its pathmnist subset, derived from 2D image classification. On the client side(2), each client device performs local training while incorporating a curvature-aware regularization term. This regularization penalizes deviations from the global model based on the estimated importance of each parameter, thereby preserving critical knowledge and improving model stability. On the server side (1), the curvature matrices and gradients collected from the clients are aggregated. The server then applies a curvature-scaled update to the global model, ensuring that parameter adjustments are inversely proportional to their estimated importance. This approach enables more efficient and robust federated optimization in non-iid settings characterised by unbalanced data distributions  [32]. A private ethereum blockchain network is utilised as a service layer to demonstrate its capability to uphold auditability, verifiability, and availability while ensuring privacy preservation in edge-FL settings.
@@ IV-A Parameter Settings
To evaluate the proposed framework under realistic conditions, we utilise three distinct benchmark datasets, each selected to inquire different capabilities of the system. The mnist dataset, a standard benchmark for image classification, includes 70,000 images (28×28 pixels), split into 60,000 training and 10,000 test samples, providing a baseline for foundational algorithm validation. For a more clinically relevant and challenging benchmark, we employ the medmnist subclass pathmnist (28×28 pixels), a more challenging alternative, comprises 35,000 training images and 8000 testing images in non-iid, enabling evaluation of model generalizability in scenarios with higher intra-class variability. Thirdly, cifar-10, a widely used dataset for object recognition, offers 60,000 color images (32×32 pixels). Together, these datasets simulate real-world edge-FL challenges, such as decentralised computation and heterogeneous data privacy requirements, while validating the framework’s ability to balance transparency, security, and efficiency in privacy-sensitive environments.
@@ IV-B Network Configuration:
This study utilizes two widely recognised neural architectures: a multi-layer perceptron (MLP) and a convolutional neural network (CNN). These models served as the foundational deep learning frameworks for training classification systems within a FL setup, simulating server client training scenarios. The experiments aimed to assess how effectively each algorithm handles non-iid data and preserves knowledge across distributed medical and non-medical imaging tasks.
For the image-based datasets, mnist and pathmnist, the CNN architecture consists of two convolutional layers, each followed by a max-pooling layer for spatial downsampling, and concludes with two fully connected layers. To accommodate the RGB input channels of the cifar-10 dataset, the architecture is modified while retaining the core structure. Both configurations employ the SGD optimizer with a learning rate of 0.001.
@@ IV-C Hardware and Software Configuration
We utilise raspberry pi3 model B+ minicomputers as edge device servers (manager nodes of side chains) and lightweight nodes are implemented using STM32F427 development boards (low-power ARM cortex M3,M4 and M7 processors), which are used for high-speed implementation of asymmetric cryptographic algorithms. ECDSA is used to generate public and private keys, and device authentication mechanisms. The STMicroelectronics x-cube-cryptolib library is utilised to implement several standard cryptographic algorithms with the ARM cortex-M series processors. System components are developed using golang, solidity to write smart contracts and deployed using remix IDE, and metamask handles concurrent transactions and interactions with third-party cloud services. A PoS consensus ensures block validation, while the gossip protocol enables fast, resilient message propagation and node synchronization. Figure 8 depicts the system’s initialised nodes.
@@
   Compute regularised loss
@@ V Performance Evaluations
To validate the effectiveness of BFEL, we conducted experiments across varying task sequence configurations using FedCurv and compared its performance against established baselines using FedAvg. The results are analysed through multiple perspectives, offering distinct insights. Code is available publicly at our github account [^1] https://github.com/AnumNawazKahloon/FedCurv.
@@ V-A Federated Simulation Results
From the epochs per round standpoint, accuracy consistently improves as the number of local training epochs increases across all settings and algorithms. This trend depicts the close alignment of local optima with global optima, making extended local training within each round advantageous when maintaining a fixed number of communication rounds. Accuracy of FedAvg increases sharply after each epoch as compared to FedCurv at the base level depicted in figure 7 and  6 as well as edge client level, which shows the minimum divergence of results after each round in FedCurv, makes it more consistent and according to the previous weights of edge client devices.
In an overall comparison of performance, FedCurv designed primarily to address non-iid data challenges in FL, surprisingly outperforms FedAvg even in uniform data settings. Notably, FedCurv often achieves superior accuracy after 100 rounds, indicating a potential need for extended training phases to reach convergence compared to FedAvg. Regarding communication efficiency, reducing the frequency of communication rounds while keeping the total number of training epochs constant yields better model performance, implying that less frequent parameter exchanges may enhance learning stability or optimization.
The experimental setup employed the SGD optimizer with an adaptive learning rate decay, reduced by a factor of three after every five epochs, along with a mini-batch size of 20 wit 20 rounds per task, and 1 epoch per round. For the MLP configuration, adjustments included a reduced mini-batch size of 10 and an initial learning rate of $1\mathrm{e}{-4}$, while client sampling fractions of 0.25 and 0.05 were applied at each round. Hyperparameters $\lambda 1$ and $\lambda 2$ were fixed at $1\mathrm{e}{-1},4\mathrm{e}{-1}]$ and 100, respectively, across all experiments. These findings collectively highlight the interplay between local training intensity, algorithmic robustness, and communication strategies in FL frameworks.
@@ V-B Layer 2 Ethereum Implementation
The proposed system implements a second layer ethereum network is implemented using a polygon sidechain structure, which operates on e proof of stake (PoS) consensus. It consists of two main components, a primary blockchain layer hosted on cloud servers and a sub-blockchain network of multiple sidechains. Sidechains are deployed on individual edge clients to facilitate localised model training in BFEL. Within each sub-blockchain, the corresponding edge client functions as a miner node, responsible for performing local model training and updating local model weights based on the global model parameters after each training epoch. Subsequently, each edge client transmits its locally trained results to the main blockchain layer, where global model parameters are aggregated.
@@
The edge clients are also tasked with collecting and packaging model training outputs (treated as transactions) into candidate blocks, which are then published to their respective private blockchain network. The main blockchain layer is operated by cloud servers configured as a consortium blockchain. Each registered user organization such as hospitals or governmental agencies within the consortium is permitted to establish its own private sidechain. Furthermore, to reduce potential latency caused by complex task assignment mechanisms, transaction packaging responsibilities in both private and consortium blockchains are delegated to designated manager nodes.
@@ V-C Security Analysis
We have conducted a performance evaluation of three different types of STM32F427 M series processors within the framework of asymmetric cryptography. We utilised the x-cube-cryptolib library to implement the ECDSA. To ascertain the statistical error of the results obtained over the number of executions, we calculated the mean, standard deviation, and standard error using the appropriate equations.
@@
This study evaluates power consumption, a key microcontroller parameter, during the execution of cryptographic algorithms. Power was measured by calculating current consumption from the voltage drop across a shunt resistor (R) using Ohm’s law. The current consumption of the processors was calculated using ohm’s law. To determine the average power consumption, ECDSA was executed for 15 runs and a comparison of power consumption is presented in figure 9(b). The average power consumption by M3 and M4 was $\pm$ 200mW, whereas M7 used an average of $\pm$ 290mW. Results indicate the superior performance of M4 cortex microcontrollers are best fit while consuming fewer resources.
@@
A deterministic wallet is used for ckd functions to determine a child’s key from a parent’s key. Using this technique, each batch of data to be encrypted in the device is given a unique secret key  [33]. The 512-bit hash is calculated according to the parent’s public key (public and private keys are 256 bits) and the desired child index. It is impossible to deduce the original parent key from the nthth-child key because of the one-way hashing used in the process. This process appears to generate random numbers due to the additions modulo n.
@@ V-D Scalability Analysis
Client based DLTs offer several offer significant benefits, but also face inherent limitations in terms of scalability limitations that restrict number of parallel processes. Nevertheless, the BFEL handles this challenge using the side chains concept, a gossip protocol, and PoS consensus. FoBSim simulation tool [^2] https://github.com/sed-szeged/FobSim is utilised to check the scalability of the proposed model. Manager nodes ranging from 5 to 500 were used to check the performance matrix of the proposed model and measure the total time required to complete the request procedure at clients level devices versus at the cloud layer.
The client measures and divides the time needed to complete a transaction into three sections: time to retrieve data (TRD), time for checking the transaction (VTR), and time for confirming the transaction (TCT). Figure 10(a) presents the results of the measurements. Based on the available resources, it is impressive that TRD requires only 34.6 milliseconds on average, VTR 36 milliseconds on average, and TCT 73.6 milliseconds on average. Additionally, it is essential to note that TCT also relies on the network, which in this experiment was adversely affected by our shared Wi-Fi’s slow response time, causing the time to be extended overall.
The gossip protocols are integrated to improve the scalability of the proposed network model and public key infrasture (PKI) is utilised for cryptographically signing model updates and transactions. Results in  10(b) demonstrates the clear reduction in total elapsed time compared to transactions processed without gossip protocol. Figure 11(a) shows that the cloud layer utilises around double the time as compared to edge clients to complete transaction requests during concurrent transactions starting from 5 to 100 transactions at a time. This architecture integrates end-to-end privacy pipelines, tamper evident logging, and granular access controls, ensuring regulatory compliance and transparency across all operational phases.
@@ VI Discussion and Analysis
In this study, we propose and develop a personalised healthcare system BFEL based on second order FEL. Second-order FL methods bring a substantial benefit for personalised training through their capability to use loss function curvature data to improve personalised training procedures during local data training. The BFEL framework integrates blockchain as a service layer to ensure data privacy, transparency and confidentiality of data owners while maintaining auditability, verifiability, availability, and robust security across FEL environments. Upon model training completion, both global model and local parameters are securely stored on the blockchain and subsequently distributed to the FL edge client devices and cloud servers in accordance with the established policy. Smart contracts are utilised to broadcast policies while providing automated confirmation and distribution operations.
To enhance personalized health monitoring and prediction[34], we employ an optimised FedCurv algorithm. It minimizes the necessary communication rounds for training on non-iid and heterogeneously distributed data across edge client devices. Performance evaluations demonstrate better performance of FedCurv and resistance capabilities than FedAvg using CNNs and MLPs in non-iid conditions. Key evaluation metrics including throughput, accuracy, privacy, and scalability analysis.
Analysis of the experimental results demonstrates that the proposed method preserves the throughtput and accuracy of the ML process while ensuring auditability and verifiability throughout the training and aggregation procedures. Privacy protection is achieved through second-order FEL using FedCurv and public key encryption ECDSA and ECIES, while latency and throughput were evaluated by measuring communication transactions on a permissioned blockchain and compared against a benchmark model FedAvg. The results highlight that BFEL outperforms the benchmark by achieving enhanced privacy, accuracy and scalability. The results underscore better stability and accuracy of the proposed framework BFEL in heterogeneous settings and non-iid environments where patient data varies widely.